OpenStudy (misssmartiez):

Which expression is equivalent to 3^2 ⋅ 3^−5?

OpenStudy (misssmartiez):

1/3^3 1/3^7 1/3^-3 1/3^-7

OpenStudy (3mar):

Do you mean \[3^2*3^{-5}\]?

OpenStudy (misssmartiez):

Yes

OpenStudy (3mar):

Ok

OpenStudy (3mar):

You and I know that we sum the power (index) in the case of multiplication terms with the same bases. Good?

OpenStudy (misssmartiez):

Yes

OpenStudy (3mar):

\[a^c*a^d=a^{c+d}\]

OpenStudy (3mar):

Great!

OpenStudy (misssmartiez):

1/3^2- (-5)?

OpenStudy (misssmartiez):

I know its either A or D

OpenStudy (3mar):

so the result will be\[3^{2+(-5)}=3^{2-5}=3^{-3}=\frac{ 1 }{ 3^3 }=\frac{ 1 }{ 27 }\]

OpenStudy (misssmartiez):

B OR D))

OpenStudy (3mar):

The first choice

OpenStudy (misssmartiez):

Oh, wait, hold on, i did division

OpenStudy (3mar):

Tell me that you got that or not.

OpenStudy (3mar):

Yes, you did. Instead do multiplication. You got the mistake yourself!